21 November 2007

RHIC Startup:

The experimental access began today a little early (8:30AM)  Both are at restricted access.   In addition, there will be an attempt to fix a broken IPM, and a swap of a power supply card at yi10-tq4.

Q89 in yellow is exceeding the current limit of 275Amos.  It is at 289A.  Last year this supply gave a lot of problems at this current.  Lowering the limit to 275A made most of the problems disappear.  The hall is cooler this year, so it may be OK now, but Don would prefer to operate at lower current.  Instead of operating at such a high current, we need to develop a new lattice that does not require such a high current in q89.  SteveT needs to work on this during the access today.   

The rebucketing done earlier in the week appears to need to be redone.  This was a "data management" error.  Now that we have rebucketed beam, it is time to turn on the stochastic cooling.  There was a lot of spillover, gold more than deuterons.  There is work to be done to get tighter bunches.

 Background rates were about 4 times as high as they need to be.  So it is clear that we need to recheck the collimation, especially if we redo the lattice. 

Gap cleaning was working yesterday, but was not turned on in this morning's stores.


Scheduling Physicist: Mike Sivertz Last updated:
Email:sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020